# Statistical Test for Data Analysis Pipeline by Selective Inference

### v1

@@ Appendix G Examples of Implementations
11 O = plp.cook_distance(X, y, 3.0)
12 X, y = plp.remove_outliers(X, y, O)
@@
14 M1 = plp.stepwise_feature_selection(X, y, 3)
15 M2 = plp.lasso(X, y, 0.08)
16 M = plp.intersection(M1, M2)
17 return plp.make_pipeline(output=M)
@@
21M, p_list = pl.inference(X, y, sigma=1.0)
Listing 3: Code example that defines the automatic pipeline construction scheme referred to as all_cv in the experiments. We can implement an instance of the tunable pipeline with fit methods simply by specifying each operation and its candidates of parameters in turn. To perform hypothesis testing after cross-validation, we can call the fit and inference method of the pipeline instance sequentially.
@@
4def option1_cv():
5 X, y = plp.make_dataset()
6 y = plp.mean_value_imputation(X, y)
@@
8 O = plp.soft_ipod(X, y, 0.02, {0.02, 0.018})
9 X, y = plp.remove_outliers(X, y, O)
@@
11 M = plp.marginal_screening(X, y, 5, {3, 5})
12 X = plp.extract_features(X, M)
@@
14 M1 = plp.stepwise_feature_selection(X, y, 3, {2, 3})
15 M2 = plp.lasso(X, y, 0.08, {0.08, 0.12})
16 M = plp.union(M1, M2)
17 return plp.make_pipeline(output=M)
@@
19def option2_cv():
20 X, y = plp.make_dataset()
21 y = plp.definite_regression_imputation(X, y)
@@
23 M = plp.marginal_screening(X, y, 5, {3, 5})
24 X = plp.extract_features(X, M)
@@
26 O = plp.cook_distance(X, y, 3.0, {2.0, 3.0})
27 X, y = plp.remove_outliers(X, y, O)
@@
29 M1 = plp.stepwise_feature_selection(X, y, 3, {2, 3})
30 M2 = plp.lasso(X, y, 0.08, {0.08, 0.12})
31 M = plp.intersection(M1, M2)
32 return plp.make_pipeline(output=M)
@@
34pls = plp.make_pipelines(option1_cv(), option2_cv())
@@
37pls.tune(X, y, n_iters=[16, 16], cv=5, random_state=0)
38M, p_list = pls.inference(X, y, sigma=1.0)

### v2

@@ 9 X = extract_features(X, M)
11 O = cook_distance(X, y, 3.0)
12 X, y = remove_outliers(X, y, O)
@@
14 M1 = stepwise_feature_selection(X, y, 3)
15 M2 = lasso(X, y, 0.08)
16 M = intersection(M1, M2)
17 return construct_pipelines(output=M)
@@ 16 M = intersection(M1, M2)
21M, p_list = manager.inference(X, y, sigma=1.0)
Listing 3: Code example that defines the automatic pipeline construction scheme referred to as all_cv in the experiments. We can create an instance of manager class which handles identically structured pipelines, each with a different hyperparameter set, simply by specifying each operation and its candidates of parameters in turn (corresponding to option1_multi and option2_multi). Manager instances can use the OR operator | to create new manager instance which handles all of the pipelines that each instance handles. To perform hypothesis testing after cross-validation, we can call the fit and inference method of the manager instance sequentially.
@@
4def option1_multi() -> PipelineManager:
5 X, y = initialize_dataset()
6 y = mean_value_imputation(X, y)
@@ 5 X, y = initialize_dataset()
8 O = soft_ipod(X, y, [0.02, 0.018])
9 X, y = remove_outliers(X, y, O)
@@
11 M = marginal_screening(X, y, [3, 5])
12 X = extract_features(X, M)
@@
14 M1 = stepwise_feature_selection(X, y, [2, 3])
15 M2 = lasso(X, y, [0.08, 0.12])
16 M = union(M1, M2)
17 return construct_pipelines(output=M)
@@ 16 M = union(M1, M2)
19def option2_multi() -> PipelineManager:
20 X, y = initialize_dataset()
21 y = definite_regression_imputation(X, y)
@@ 20 X, y = initialize_dataset()
23 M = marginal_screening(X, y, [3, 5])
24 X = extract_features(X, M)
@@
26 O = cook_distance(X, y, [2.0, 3.0])
27 X, y = remove_outliers(X, y, O)
@@
29 M1 = stepwise_feature_selection(X, y, [2, 3])
30 M2 = lasso(X, y, [0.08, 0.12])
31 M = intersection(M1, M2)
32 return construct_pipelines(output=M)
@@ 31 M = intersection(M1, M2)
34manager = option1_multi() | option2_multi()
@@
37manager.tune(X, y, num_folds=2)
38M, p_list = manager.inference(X, y, sigma=1.0)